OpenStudy (anonymous):

220e^(1.323t)=500e^(1.118t) solve for t

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

so when u divide by 500 u get: 0.44e^(1.323t) = e^(1.118t)

OpenStudy (mertsj):

Here is a better approach: 1. divide both sides by e^(1.118t)

OpenStudy (anonymous):

that is: 220e^(1.323t)/e^9.118)= 500

OpenStudy (mertsj):

\[220e ^{.205t}=500\]

OpenStudy (anonymous):

do u subtract 1.323-1.118?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Now divide both sides by 220

OpenStudy (anonymous):

yup & iget 2.272727......

OpenStudy (anonymous):

or should i just do: ln(500/220) / 0.205?

OpenStudy (mertsj):

\[e ^{.205t}=2.272727...\]

OpenStudy (mertsj):

Now take the ln of both sides.

OpenStudy (anonymous):

so i got t= 4.005 ?

OpenStudy (mertsj):

\[.205t=.82098\]

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

thank u! :)

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

i have 1 more question. can i start a new question & tag u?